The perimeter of a rectangular backyard is 6x +6 yards. If the width is x yards , find a binomial that represents the length.
Answer by nyc_function(2741)

Let p = perimeter
Let L = length
Let w = width
p = 2L + 2w
6x + 6 = 2L + 2x
6x + 6 - 2x = 2L
4x + 6 = 2L
(4x + 6)/2 = L
2x + 3 = L